OpenStudy (anonymous):

A plant 8 inches tall is placed by a window. If it grows 3 inches a day, how many days will it take 26 inches tall?

OpenStudy (anonymous):

Concept = rates 3 inches per day is the rate here, and we start out at 8 so...

OpenStudy (btaylor):

26=8+3x 18=3x x=6 days

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[26 \text{inches} = ( 3 \ \frac{\text{inches}}{\cancel{\text{day}}})(x \ \cancel{\text{days}}) + (8 \ \text{inches})\] Don't forget the units! :-)